# Proactive Agents for Multi-Turn Text-to-Image Generation Under Uncertainty

### v1

@@ 4 Proactive T2I agent design
We provide high-level principles and design that guide our agent how to behave and interact with users to generate desired images from text through multi-turn interactions. The goal of the agent is to generate images that match the user’s intended image as closely as possible with minimal back-and-forth, particularly in cases with underspecified prompts and the agent needs to gather information proactively. This requires a decision strategy on information gathering to trade off between the cost of interactions and the quality of generated images. The formal problem definition can be found in §C.
We equip the agent with the ability to gather information in two ways: ask clarification questions (§4.1) and express its uncertainty and understanding in a way that users can edit (§4.2). Once a piece of information is collected from a user, the agent also need to update its questions and uncertainty (§4.3). To enable all these agent behaviors, we need to situate the agent in an interface to effectively communicate with users (§G). In the following, we introduce the design of the above components under the interface, to ensure information efficiency for T2I generation.
4.1 What kind of questions should be asked?
We explain considerations in question asking and examples of strategies in this section.
4.1.1 Principles
@@
We identify the following principles for an agent to ask the user questions about the underspecified prompt and their intended image: (i) Relevance: The question should be based on the user prompt. (ii) Uncertainty Reduction: The question should aim to reduce the agent’s uncertainty about the attributes and contents of the image, the objects, the spatial layout, and the style. (iii) Easy-to-Answer: The question should be as concise and direct as possible to ensure it is not too difficult for the user to answer. (iv) No Redundancy: The question should not collect information present in the history of interactions with the user. The Relevance and No Redundancy principles are self-explanatory, we detail the other two principles below.
@@
The Uncertainty Reduction principle aims to let agent elicit information about various characteristics of the desired image, which the agent is unsure of.
@@
First, the agent needs to know what characteristics of images are important. Some examples include: (i) Attributes of the subjects, such as breed, size, or color, with questions like What kind of rabbit? What color is the cat?; (ii) Spatial relationships between the subjects, such as proximity and relative position (Are the rabbit and cat close to each other? Are they facing each other?); (iii) Background information, such as location, style and time of day (Are they in a park or at home?); and (iv) Implicit entities that might not be explicitly mentioned in the initial prompt but are relevant to the user’s vision (Are there any other animals or people present?).
@@
Second, the agent needs to know its own uncertainty about those characteristics. In the agent’s belief, the uncertainty is explicit. One strategy is to form questions about the image characteristics that the agent is most uncertain about. We discuss more in §E.2.
@@
The Easy-to-Answer principle aims to reduce users’ effort to respond to questions. One way is to have the agent provide some answer options, where options are what the agent believes likely to appear. E.g., What color is the cat? (a) Black (b) Brown (c) Orange (d) Other (please specify).
4.1.2 Examples of question-asking strategies
Given the agent belief constructed from the user prompt (more details in §4.2), several basic approaches can be employed following the above principles. We construct simple agents with the following strategies, which are implemented and used in our experiments.
  • Ag1 (§E.5): Rule-based question generation, which leverages predefined rules or heuristics to identify salient attributes, entities, or relationships that require clarification. For example, an LLM could be used to estimate the importance and likelihood of different components within the belief, and a heuristic could be applied to prioritize the most crucial elements for questioning.
  • Ag2 (§E.6): Belief-guided question generation, which involves using natural language to represent the current understanding encapsulated in the belief. This representation, along with the conversation history, is provided as input to an LLM, guiding it to generate clarification questions.
  • Ag3 (§E.7): Direct question generation, which write the above question-asking principles in a prompt for an LLM to generate a question.
4.2 Interacting with the user based on agent beliefs
The Uncertainty Reduction principle inspires the usage of belief graphs for the agent to directly express uncertainty, in addition to reflecting uncertainty through questions. Instead of using hardcoded symbols in classic belief representations (Fikes and Nilsson, 1971) described in §3, we employ LLMs to generate names and values for entities, attributes and relations. As a result, this belief construction method can generalize across any prompts. Algorithm 1 summarizes how an agent parses a prompt to a belief graph and allows user interaction [^3] The clarification question part of the interaction is omitted for simplicity. All agents in §4.1.2 use the same kind of belief graphs.
Algorithm 1 Belief Parsing and interaction
1:  Input: Initial Prompt (IP)
2:  Initialization: Merged Prompt (MP) $\leftarrow$ IP
3:  for $turn\leftarrow 1$ to $max\_turn$ do
4:     Parse entities from MP (E.8)
5:     Parse entity attributes and relations from entities and MP (E.9, E.10)
6:     Display belief graph, and collect interaction feedback (F)
7:     Update MP: MP $\leftarrow$ MP + F (E.12)
8:  end for
Entities.  In addition to (a) entities mentioned in the user prompt, a belief graph also includes (b) implicit entities not mentioned in the prompt but likely to appear, e.g., pet owner in the context of a pet-related scene; and (c) background entities, such as image style, time of day, location, which play important roles in constructing the image.
Attributes and relations.  While the prompt might mention some attributes of a certain entity, they are not enough to describe the exact details of that entity. Hence the agent have to imagine the relevant attributes for each entity, and construct a list of possible values along with their associated probabilities (e.g., the color attribute for the cat entity might have values like black, white, gray with corresponding probabilities). Similarly the agent may have to imagine the possible relations between entities, e.g., spatial relation between rabbit and cat might include values like close, far, touching.
Importance scores.  While the agent can be uncertain about many aspects of the user’s intended image, some are more important than others. E.g., for prompt “a rabbit and a cat”, the agent might be very uncertain about the exact color of a carpet that might appear in the image, but rabbit and cat are more important than the carpet. We enable agents to estimate an importance score for each entity, attribute and relation.
Extracting beliefs and enabling interactions.  A simple idea is to use a large language model (LLM) via in-context learning. §E.1 details how an LLM may analyze the user prompt to identify entities, their attributes, and the relations between them, effectively translating the natural language input into a structured representation within the belief. Once the belief is extracted, a user can edit the belief to adjust uncertainty levels, confirm existence of entities etc, as shown in Figure 1.
4.3 Transition
The agent belief undergoes a transition whenever the agent receives new information through user feedback, either user answers from the agent question or user interactions with the graph-based belief interface (Figure 1). This transition process integrates information from the initial user prompt, the conversation history, interaction and the previous belief to generate an updated belief of the user’s desired image. We use a simple approach: Generate a comprehensive prompt that summarizes all interactions and information gathered thus far. This merged prompt is then used to re-generate the belief, effectively incorporating the new information into a refreshed representation. The implementation details can be found in §E.3.
@@ 5.1 Automatic evaluation
We simulate the user-agent conversation using self-play (Shah et al., 2018) between two LLMs. The conversation starts with an arbitrarily chosen image to represent the goal image from a T2I model that the user has in mind [^4] This assumption only applies to the experiments. In practice, users don’t necessarily have an image in mind, but they can get inspirations from the belief graphs and questions.. Along with this ground truth image, a user has a detailed prompt (i.e., the ground truth) in mind that describes the image in high-detail. We use the algorithm similar to Ag2 (detailed in §E.4) to simulate the user, where the questions are answered based on the ground truth prompt and the belief graph generated from the ground truth prompt. We run the agent-user conversation for a total of 15 turns [^5] While 15 turns is a suggested approximation of interaction time, accounting for varying difficulty between images, any number of turns can be used with this evaluation approach. and compute different metrics at the end of each turn. More details of the simulated user can be found in the appendix, including the prompts provided to the LLM when simulating the user are provided. Figure 2 part b shows the multi-turn set up that we use in our results.
Figure 2: a) Each column displays the output of an agent after 15 turns - the right most column shows target image, which belongs to DesignBench. b) A visualization of the multi-turn set up in the experiments. These are real generated outputs and simulated user outputs at turns 3, 10 and 15.
@@ 5.1.1 Setups for agents and baseline
Baselines.
We use a standard T2I model as a baseline, which directly generates an image based on a prompt without asking any questions. We refer to this baseline as ‘T2I’.
Agents.
We use Ag1, Ag2 and Ag3 with question-asking strategies introduced in §4.1.2. The creation and updates to the belief graph (§4.2), as well as transitions to prompt (§4.3) are consistent among all multi-turn agents. Further implementation details of each agent can be found in §E.
Model Selection.
In this work we use an off-the shelve Text-to-Image (T2I) model and a Multi-Modal Large Language (MLLM) model and build the different components of our agent on top of these models. We keep these models consistent across all agents for fair comparison. We implement the agent on top of the Gemini 1.5 (Gemini Team Google, 2024) using the default temperature and a 32K context length. The in-context examples and the exact prompt used at each step of the agent pipeline is detailed in §E.8 - §E.15. More agent implementation details are provided in §E. For T2I generation, we use Imagen 3 (Baldridge et al., 2024) across all baselines given it’s recency and prompt-following capabilities. We used both the models served publically using the Vertex API [^6] https://cloud.google.com/vertex-ai.
@@
Our multi-turn agents aim to facilitate the generation of complex images, a process that often requires users to iteratively refine text-to-image (T2I) prompts until the generated image aligns with their mental picture. To evaluate these agents, we curate datasets comprising complex scenes involving multiple subjects, interactions, backgrounds, and styles. Each dataset consists of tuples: $(\mathbf{I},p_{0},c,b_{gt})$, where $\mathbf{I}$ represents the target image, $p_{0}$ is an initial (basic) prompt describing only the primary elements of the scene, $c$ is a ground truth caption providing a detailed description of $\mathbf{I}$, including spatial layout, background elements, and style, and $b_{gt}$ is the ground truth belief graph constructed via parsing $c$. The initial prompt $p_{0}$ is intentionally less detailed than $c$ to necessitate multi-turn refinement. This framework allows us to assess the agent’s ability to guide the user towards the target image $\mathbf{I}$ starting from a simplified prompt.
Existing image-caption datasets primarily focus on simple scenes (Deng et al., 2009; Krizhevsky et al., 2009; Deng, 2012) or focus on very specific categories (Liu et al., 2016; Liao et al., 2022). With the aim for complex realistic images for testing the robustness of the agents, we evaluate over the validation split of the Coco-Captions dataset (Chen et al., 2015). Five independent human generated captions are provided for each image in the dataset. These captions are often short and describe the basic elements contained in the image and the interactions between objects or persons in the image. We therefore select the shortest of the five human-generated captions and use this as a starting prompt $p_{0}$. We then use Gemini 1.5 Pro to expand the starting prompt by adding more details of the attributes of the entities in the image as well as the style and image composition which results in the ground truth caption. We also use the ImageInWords (Garg et al., 2024) dataset which takes a diverse set of realstic and cartoon images and has human annotators create dense detailed captions that describe attribute and relationships between objects in the image. In ImageInWords evaluations we use the long human annotation as the ground truth caption.
While COCO-Captions and ImageInWords provide complex, real-world images across diverse backgrounds, it lacks the artistic or non-photorealistic imagery often desired by designers and artists seeking to generate content outside the distribution of typical training data. To better evaluate our target for flexible use cases such as by artists, we introduce DesignBench, a novel dataset comprising 30 scenes specifically designed for this purpose. Each scene follows the $(\mathbf{I},p_{0},c,b_{gt})$ format described earlier. DesignBench includes a mix of cartoon graphics, photorealistic yet improbable scenes, and artistic photographic images. Examples from DesignBench and a comparison with COCO-Captions are provided in the Appendix.
@@ 5.1.3 Metrics
The outputs produced by the agent include a final generated image, a final caption and a final belief graph. We evaluate the agents across these modalities and evaluate their alignment to the ground truth image $\mathbf{I}$, caption $c$ and belief $b_{gt}$, using the following metrics.
Text-Text Similarity: We use 2 metrics for comparing the ground truth caption and the generated caption: 1) T2T – embedding-similarity computed using Gemini 1.5 Pro [^7] Text embeddings are obtained from Embeddings API: https://ai.google.dev/gemini-api/docs/embeddings. and 2) DSG (Cho et al., 2024) adapted to parse text prompts into Davidsonian scene graph using the released code.
@@
Text-Image Similarity: We compare the ground truth prompt with the generated image (T2I) using VQAScore (Lin et al., 2024). We use the author released implementation of the metric and use Gemini 1.5 Pro as the underlying MLLM.
@@
Negative log likelihood (NLL): We construct the ground truth state of the image in the form of a belief graph but with no uncertainty. We then approximately compute the NLL of the ground truth state given the belief of the agent at each turn, by assuming the independence of all entities, attributes and relations, and summing their log probabilities [^8] This approximation does not account for potential similarities in the names of entities or attributes. This could lead to approximation errors if, for example, the model confuses ”Persian cat” with ”Siamese cat” due to their similar names. Addressing this limitation would require incorporating semantic similarity measures into the NLL computation..
5.2 Results from automated evaluation
@@
The results from the automatic evaluations in Table 1 show the $\mathbf{I}$, $c$ and $b_{gt}$ against each agents final generated image, text and state. All show the mean and standard deviation of the similarity metric at the final agent state. The blue row shows the baseline method which performs no updates to the prompt and instead applies the T2I model to the first prompt. Therefore this baseline represents the lower bound performance.
To add quantitative validity to the ground truth caption generation we perform Text to Image (VQA) Similarity between the ground truth caption and the ground truth over all images in the DesignBench dataset. The mean T2I VQA similarity between the ground truth caption and ground truth image is 0.99999985 with a median 1.0, and standard deviation of 4.5e-07. The mean is extremely close to 1 as expected of an accurate and well formed caption. These numbers can be compared to the T2I column of Table 1 to observe the delta between the ground truth caption and generated captions.
@@
The results in Table 1 show that significant gains in performance come from using proactive multi-turn agents. The blue row shows the simplest baseline which directly uses a T2I model and performs no updates to the initial prompt $p_{0}$. We see that all of the multi-turn agents far exceed the baseline T2I model on both datasets and all metrics. Ag3 (the LLM agent that does not explicitly utilize the belief graph to generate questions) show superior performance across all metrics.
The plots in Figure 3 show the T2T, I2I, T2I and NLL metrics, averaged across all images in the ImageInWords dataset, per turn for 15 turns. We see that the multi-turn agents all improve in every metric as they increase the number of interactions. Interestingly we see the T2T and the T2I VQA similarity metric seems to plateau or decrease after about 10 interactions, while the I2I scores continue to increase. The NLL metric shows large performance gains of the Ag3 agent in comparison to all other methods. The plots in Figure 12 shows the T2T DSG metrics.
@@
Table 1: Automatic evaluation results on Coco-Captions, ImageInWords, and DesignBench. Our agents show large performance gains in all metrics over a standard T2I model alone.
5.3 Analysis of quantitative results
The evaluations on the COCO-captions, ImageInWords, DesignBench datasets show similar results and highlight the same patterns across the different agents.
Multi-Turn agents show clear advantage: The immediate take away is the baseline which does not use multi-turn interaction and instead passes in the original prompt into the T2I model performs worse than the multi-turn agents on all metrics on both datasets. This confirms our hypothesis that the current T2I agents often produce less desirable images given ambiguity in prompts. In Figure 2 we see real outputs of the multi-turn set up with the Ag3 agent.
LLMs being a part of agents play a significant role: The best performers (Ag2 and Ag3) both query and LLM to provide a question to ask the user based on contextual information such as the belief graph and conversation history. They query the LLM to construct a concise and clear question but don’t impose further constraints on the question construction. Ag1 provides a programatic template for how the LLM should construct the question based on its belief graph and does not provide any conversation history information. Examples of dialogs and the generated questions produced by the three agents can be found in the Appendix in Figure 6. This figure demonstrates that the templated question creation leads to extremely specific questions that often gather minimal information in return. This is an intrinsic limitation of hard coded question selection strategy but also can be an issue of the heuristic scores we defined for question selection in Ag1. In contrast, Ag2 and Ag3 generate questions that are more open-ended thus allowing the user to provide more nuanced details which in consequence enhance the agent’s image knowledge.

### v2

@@ 5 Proactive T2I Agent Design
We provide high-level principles and design that guide our agent how to behave and interact with users to generate desired images from text through multi-turn interactions. The goal of the agent is to generate images that match the user’s intended image as closely as possible with minimal back-and-forth, particularly in cases with underspecified prompts and the agent needs to gather information proactively. This requires a decision strategy on information gathering to trade off between the cost of interactions and the quality of generated images. The formal problem definition can be found in §B. Algorithm 1 summarizes how the agent constructs an initial belief graph from user prompt, takes actions (including asking questions and presenting belief graphs), incorporates user feedback during interaction, and updates the belief graph and prompt.
Algorithm 1 Belief parsing and interaction
1:  Input: Initial Prompt ($p$)
2:  Initialization:
3:  Parse entities from prompt $p$ (D.5)
4:  Parse entity attributes and relations from entities and prompt $p$ (D.6, D.7)
5:  Construct initial belief graph $b$
6:  for $turn\leftarrow 1$ to $max\_turn$ do
7:     Take action $a\leftarrow\pi(b,p)$
8:     Observe user feedback $o$
9:     Update $b,p\leftarrow\tau(b,p,o)$ (D.9)
10:  end for
5.1 Belief Graph Transitions
The agent’s belief graph gets updated when the agent receives new information through user feedback. This feedback can come from agent-user dialog or through a user direct manipulation of the belief graph on the agent interface (Figure 1). The transition process integrates information from the initial user prompt, the conversation history, interaction and the previous belief. In an agent prototype, we use a simple approach to implement the transition: Generate a comprehensive prompt that summarizes all interactions and information gathered thus far. This merged prompt is then used to re-generate the belief, effectively incorporating the new information into a refreshed representation. The implementation details can be found in §D.4.
5.2 Asking Questions to Gather Information
@@
We identify the following principles for an agent to ask the user questions about the underspecified prompt and their intended image: (i) No Redundancy: The question should not collect information present in the history of interactions with the user. (ii) Uncertainty Reduction: The question should aim to reduce the agent’s overall uncertainty about what contents should be included in the generated image, e.g., entities, attributes, spatial layout, style. (iii) Relevance: The question should be based on the user prompt. (iv) Easy-to-Answer: The question should be as concise and direct as possible to ensure it is not too difficult for the user to answer. The No Redundancy and Relevance principles are self-explanatory, we detail the other principles below.
@@
The Uncertainty Reduction principle aims to let the agent elicit information about various characteristics of the desired image, which the agent is unsure of.
@@
First, the agent needs to know what characteristics of images are important. For the “Image of a rabbit near a cat” prompt, some examples include: (i) Attributes of the subjects, such as breed, size, or color, with questions like What kind of rabbit? What color is the cat?; (ii) Spatial relationships between the subjects, such as proximity and relative position (Are the rabbit and cat close to each other? Are they facing each other?); (iii) Background information, such as location, style and time of day (Are they in a park or at home?); and (iv) Implicit entities that might not be explicitly mentioned in the initial prompt but are relevant to the user’s vision (Are there any other animals or people present?).
@@
Second, the agent needs to know its own uncertainty about those characteristics. In the belief graph of the agent, the uncertainty is explicit. One strategy is to form questions about the image characteristics that the agent is most uncertain about. We can define an acquisition function that takes into account the entropy over entities, attributes, relations, as well as the importance scores. We discuss more in §D.2.
@@
The Easy-to-Answer principle focuses on decreasing the cognitive load on users when they are asked to answer questions. A common strategy is for the agent to offer pre-defined answer choices. These choices are typically the ones the agent deems most probable. For example, What color is the cat? (a) Black (b) Brown (c) Orange (d) Other (please specify).
6 Implemented Agent Prototypes
We propose and experiment with three agent prototypes (Ag1, Ag2, and Ag3). All agents use the same belief graph construction approach in Algorithm 1. They differ in how they address the four question-asking principles (detailed in §5.2), specifically through their decision strategy $\pi$.
6.1 Ag1: Heuristic Score Agent
Ag1 asks questions about an entity’s existence, attribute values, or relation types when these elements are both highly important and uncertain. It calculates a heuristic score in Equation 1 by multiplying importance scores and entropy of every belief graph element, then selects the highest-scoring question. We then prompt an LLM (§D.12) to craft easy-to-answer, multiple-choice questions like, “What color of the rabbit do you have in mind? a. black, b. white, c. brown, d. unknown, or something else?”
For belief graph transition, Ag1 re-generates the belief graph based on the merged prompt ($p$ in Line 9 of Algorithm 1), and applies the post-processing logic outlined in §D.4 to ensure consistency and prevent redundancy.
Figure 4 shows Ag1 excels at constructing easy-to-answer questions. However, its hardcoded post-processing and heuristic function doesn’t consistently prevent redundancy, largely due to potential LLM errors in parsing beliefs from prompts. For example in Figure 5, Ag1 asks whether “cake” exists, even when it was mentioned in the initial prompt. This can happen if the LLM assigns “cake” a less than 100% existence probability and a high importance score.
6.2 Ag2: Belief-prompted Agent
Ag2 employs a belief-based $AICQ_{B}$ prompt in §D.11: instead of using a heuristic score like Ag1, it directly prompts an LLM to produce a question. The LLM generates each question using the belief graph, merged prompt, and conversation history. This comprehensive input allows it to craft highly relevant, informative and non-redundant queries. The LLM is also guided with specific instructions on how to construct ideal questions, such as “The question should be as concise and direct as possible,” to make questions less difficult to answer, however they are not limited to multiple choice. Questions are often of the structure “What color is the rabbit in the image?”
Ag2 does not require post-processing logic for the belief transition, since the LLM is expected to use the conversation history to judge redundancy. Our experiments in §7.2 support this hypothesis, showing that Ag2 tends to achieve better performance than Ag1 with fewer turns of questions. However, Ag2 generates fewer easy-to-answer questions than Ag1, as shown in Figure 4.
6.3 Ag3: Principle-prompted Agent
The primary distinction between Ag2 and Ag3 is the information they are prompted with. While Ag2 leverages a belief graph and merged prompt, Ag3 uses the $AICQ_{base}$ prompt (§D.10), which exclusively relies on the conversation history and the question-asking principles from §5.2.222We tried prompting an LLM with only history or simple instructions like “ask a question”, but the quality of the clarification questions was noticeably worse than prompting with the principles.
Ag3 tends to ask more open-ended questions such as “What else is present in the image?” and more complex questions like “Is the cake on a plate or a stand, and what is its color and shape?” Figure 4, Figure 3 and Table 1 show that these types of questions are more difficult for users to answer but gain more information about user intent.
@@ 7.1 Automatic Evaluation
We simulate user-agent conversations using LLM self-play in Algorithm 2, where the conversation begins with a goal image and a detailed ground truth prompt describing it. We use the algorithm similar to Ag2 to simulate the user, answering questions based on the ground truth prompt and the belief graph generated from the ground truth prompt. More details of the simulated user can be found in §D.3. We run each agent-user conversation for 15 turns and compute various metrics at the end of each turn. Figure 2 part b shows the multi-turn set up that we use in our results.
Figure 2: a) Each column displays the output of an agent after 15 turns - the right most column shows target image, which belongs to DesignBench. b) A visualization of the multi-turn evaluation set up in the experiments. These are real generated outputs and simulated user outputs at turns 3, 10 and 15.
@@ 7.1.1 Setups for agents and baseline
Baselines.  We use a standard T2I model as a baseline, which directly generates an image based on a prompt without asking any questions. We refer to this baseline as ‘T2I’. Agents.  We use Ag1, Ag2 and Ag3 with question-asking strategies introduced in §6. Further implementation details of each agent can be found in §D.
Model Selection. We implement the agent belief parsing and interaction in Algorithm 1 on top of the Gemini 1.5 (Gemini Team Google, 2024) using the default temperature and a 32K context length. For T2I generation, we use Imagen 3 (Baldridge et al., 2024) across all baselines given its recency and prompt-following capabilities. We used both the models served publicly using the Vertex API. [^3] https://cloud.google.com/vertex-ai
@@
The evaluation data used consists of tuples: $(\mathbf{I},p_{0},c,b_{gt})$, where $\mathbf{I}$ represents the target image, $p_{0}$ is an initial prompt describing only the primary elements of the scene, $c$ is a ground truth caption providing a detailed description of $\mathbf{I}$, including spatial layout, background elements, and style, and $b_{gt}$ is the ground truth belief graph constructed via parsing $c$. The initial prompt $p_{0}$ is intentionally less detailed than $c$ to necessitate multi-turn refinement. The evaluation framework we propose directly assesses the agent’s ability to guide the user towards the target image $\mathbf{I}$ starting from a $p_{0}$.
Existing image-caption datasets primarily focus on simple scenes (Deng et al., 2009; Krizhevsky et al., 2009; Deng, 2012) or focus on very specific categories (Liu et al., 2016; Liao et al., 2022). With the aim for complex realistic images for testing the robustness of the agents, we evaluate over the Coco-Captions dataset validation split (Chen et al., 2015) and ImageInWords (Garg et al., 2024) dataset. ImageInWords dataset contains a diverse set of realstic and cartoon images and has human annotators create dense detailed captions that describe attributes and relationships between objects in the image. In ImageInWords evaluations we use the long human annotation as the ground truth caption and auto-summarize the long caption with Gemini 1.5 Pro to create starting prompt $p_{0}$. For each image in Coco-Captions, the dataset provides multiple human generated captions which are short and describe the basic elements and the interactions in the image. We select the shortest of captions per image as the starting prompt $p_{0}$. We obtain the ground truth caption by expanding $p_{0}$ via prompting Gemini 1.5 Pro with ground truth image and asking it to add more details of the attributes of the entities in the image as well as the style and image composition.
COCO-Captions and ImageInWords datasets lack artistic and non-photorealistic imagery often desired by designers and artists. To better evaluate our target for flexible use cases such as by artists, we introduce and release DesignBench, a novel dataset comprising 30 scenes specifically designed for this purpose. Each scene follows the $(\mathbf{I},p_{0},c,b_{gt})$ format described earlier. DesignBench includes a mix of human generated cartoon graphics, photorealistic yet improbable scenes, and artistic photographic images. $p_{0}$ and $c$ in DesignBench are generated via prompting Gemini 1.5 Pro with ground truth image.
We validate the quality of the Gemini 1.5 Pro generated captions, $c$, by performing Text to Image (VQA) Similarity between $c$ and the $\mathbf{I}$. For DesignBench the mean T2I VQA similarity between the ground truth caption and ground truth image is 0.999 with a median 1.0, and standard deviation of 4.5e-07, as expected of an accurate and well formed caption.
@@ 7.1.3 Metrics
Each agent outputs included a generated image, a final caption and a final belief graph. We evaluate each modality via their alignment to the ground truth image $\mathbf{I}$, caption $c$ and belief $b_{gt}$, using the following metrics.
Text-Text Similarity: We use the following metrics to compare the ground truth caption and the merged prompt from the agent: 1) T2T – embedding-similarity computed using Gemini 1.5 Pro [^4] Text embeddings are obtained from Embeddings API: https://ai.google.dev/gemini-api/docs/embeddings. and 2) DSG (Cho et al., 2024) adapted to parse text prompts into Davidsonian scene graph using the released code.
@@
Text-Image Similarity: We compare the ground truth prompt with the generated image (T2I) using VQAScore (Lin et al., 2024). We use the author released implementation of the metric and use Gemini 1.5 Pro as the underlying multi-modal large language model.
@@
Negative log likelihood (NLL): We construct the ground truth state of the image in the form of a belief graph but with no uncertainty. We then approximately compute the NLL of the ground truth state given the belief of the agent at each turn, by assuming the independence of all entities, attributes and relations, and summing their log probabilities. [^5] This approximation does not account for potential similarities in the names of entities or attributes. This could lead to approximation errors if, for example, the model confuses ”Persian cat” with ”Siamese cat” due to their similar names. Addressing this limitation would require incorporating semantic similarity measures into the NLL computation.
7.2 Automatic Metrics and Qualitative Results
The automatic evaluations across COCO-captions, ImageInWords, DesignBench datasets show similar results and highlight the same patterns across the different agents.
@@ 7.2 Automatic Metrics and Qualitative Results
Results in Table 1 show the $\mathbf{I}$, $c$ and $b_{gt}$ against each agent’s final generated image, merged prompt and belief graph. All columns except ‘human rating’ show the mean and standard deviation of the similarity metric at the final turn. The blue row shows the baseline T2I method which generates an image directly from the starting prompt $p_{0}$.
@@
Multi-Turn agents show clear advantage: The results in Table 1 show that significant gains in performance come from using proactive multi-turn agents. The blue row shows the simplest baseline which directly uses a T2I model and performs no updates to the initial prompt $p_{0}$. We see that all of the multi-turn agents far exceed the baseline T2I model on both datasets and all metrics. Ag3 (the LLM agent that does not explicitly utilize the belief graph to generate questions) shows superior performance across all metrics. This confirms that the current T2I agents often produce less desirable images given ambiguity in prompts. In Figure 2 we see real outputs of the multi-turn set up with the Ag3 agent. Additionally, we see that the multi-turn agents generally improve in most metrics as they increase the number of interactions. The plots in Figure 3 show the T2T, I2I, T2I and NLL metrics, averaged across all images in the ImageInWords dataset, per turn for 15 turns. Interestingly we see the T2T and the T2I VQA similarity metric seems to plateau or decrease after about 10 interactions, while the I2I scores continue to increase. The NLL metric shows large performance gains of the Ag3 agent in comparison to all other methods. The plots in Figure 11 show the T2T DSG metrics.
LLMs play a significant role: The best performers (Ag2 and Ag3) both query the LLM to generate a question based on contextual information such as the belief graph and conversation history. They query the LLM to construct a concise and clear question but don’t impose further constraints on the question construction. Ag1 provides a programmatic template for how the LLM should construct the question based on its belief graph and does not provide any conversation history information. Examples of dialogs and the generated questions produced by the three agents can be found in the Appendix in Figure 5. This figure demonstrates that the templated question creation sometimes leads to questions that gather minimal information in return. This can be an intrinsic limitation of hard coded question selection strategy but also can be an issue of the heuristic scores we defined for question selection in Ag1. In contrast, Ag2 and Ag3 generate questions that are more open-ended thus allowing the user to provide more nuanced details which in consequence enhance the agent’s image knowledge.
@@
Table 1: Automatic evaluation results on Coco-Captions, ImageInWords, and DesignBench. Our agents show large performance gains in all metrics over a standard T2I model alone. Human Rating displays the % of times an image was ranked as closest to the ground truth image. All other columns correspond to automatic evaluations using metrics described in 7.1.3.
7.3 Human Studies on Generated Images and Dialogues
We conducted human studies to verify the automatic evaluations of the agents. See detailed design in Figure 19, Figure 20 and Figure 21. Information on human user recruitment and consent can be found in §H.
Table 1 shows the human rater evaluations on the generated images in comparison to the ground truth image. Participants are asked to rank the images produced by the three proposed multi-turn agents and a single-turn T2I model in terms of content and style against a ground truth image for which the original prompt was derived and the answers to the agent’s questions were derived. Approximately 550 image-dialog pairs per agent are rated using 3 human raters. The generated unlabeled images were presented in a random order and the human rater was tasked with ranking the images from best to worst in terms of content correctness. The results are in Table 1, under the column Human Rating, showing that agentic systems are selected as the best generated image over the single-turn T2I in 80%+ of cases for both content and style. Human ratings of the image directly correspond to and validate the I2I (DINO) results.
Figure 4 shows the human study results on the generated agent-simulated user dialogues. (1) Per question raters are asked to mark issues each agent question contains that could pose a disturbance to the user, i.e., ‘question is too long and complex’. Approximately 8k questions per agent are rated. The results in Figure 4 (left), show agentic systems have issues with their questions in 14% or less cases. Results show Ag2 and Ag3 suffer from too long questions while Ag1 suffers from questions that do not gain new information. (2) Raters are asked to evaluate the correspondence of each image to the agent-user dialog and original prompt. Approximately 1.5k image-dialog pairs are rated using 3 human raters. Results in Figure 4 (right) show that for all of our agents, more than 96% of the 1.5k image-dialog pairs are rated as very close or fairly close with some differences. This high rating shows the viability of the T2I model employed by the agents, as well as the agents’ ability to combine the dialogue into a coherent prompt to feed the T2I model.
Figure 4: Human study results on the dialogues between agents and simulated users. Left: The distribution of issues in agent questions - per agent, as determined by human raters. Right: The rating of how well the final generated image corresponds to the user prompt and dialogue.
7.4 Human Studies on the Agent Interface
We performed a human survey for understanding user frustrations and validation. We gathered data from 143 participants who all identified to be regular T2I users (at least once a month). Participants were presented with four hypothesized frustrations (prompt misinterpretation, many iterations, inconsistent generations, incorrect assumptions) and three potential mitigating features (clarifications, entity graph, relationship graph; more details in §G).
Table 4 in Appendix confirms the prevalence of hypothesized frustrations amongst users, with 83% experiencing occasional, frequent, or very frequent frustration due to prompt iterations, followed by 70% for misinterpretations, 71% for inconsistent generations, and 60% experiencing frustration due to incorrect assumptions. Most acutely 55% of participants reported frequent or very frequent frustration due to the prompt iteration frequency necessary. In Table 2, we report the mitigation features that are likely to help. Clarifications were reported as having the highest likelihood to help current workflows (91% could / likely / very likely to be helpful), followed by entity graphs (88% could / likely / very likely to be helpful) and relationship graphs (86% could / likely / very likely to be helpful). Clarifications were expected to deliver value immediately / very soon by 58%.
These findings highlight a strong user desire for, and a high likelihood of success for, features that make T2I generation more efficient and user-friendly by reducing iterations and mitigating misinterpretations. Full explanations of the hypothesized frustrations, mitigation and responses splits are in §G. All respondents were compensated for their time as per market rates, and were recruited by our vendor to ensure diversity across age, gender, and T2I usage in terms of models, frequency and purpose (work and non work).